In an optical fiber, the core material has a refractive index 1.6 and refractive index of cladding material is 1.3. calculate the value of acceptance angle.
1
Expert's answer
2020-04-28T09:33:18-0400

acceptance angle = "\sin^{-1}(n_1^2-n_2^2)" ="\sin^{-1}(2.56-1.69)=\sin^{-1}(0.87)=60.5^o"